### v1

# Probing Hawking Temperature Threshold via Quantum Depletion in Bose-Einstein Condensate

Abstract
We investigate the correlation between quantum depletion and Hawking temperature in a ring-shaped Bose-Einstein condensate featuring an analog black hole–white hole horizon pair, using the Bogoliubov approach. The presence of horizons is found to enhance the quantum depletion compared to horizon-free configurations, indicating a correlation between depletion and horizon dynamics. Via tuning the Hawking temperature, we observe its effect on the depletion profile. Our results show that depletion increases with Hawking temperature, and beyond a certain threshold, backreaction effects emerge, challenging the validity of the Bogoliubov approximation. We identify a viable parameter regime where the system remains both theoretically controlled and experimentally accessible, offering insight into horizon-induced quantum fluctuations, with implications for future studies of backreaction.
Black holes represent some of the most enigmatic objects in the universe, providing an exceptional framework for examining how general relativity and quantum theory intersect. The pioneering research by Bekenstein and Hawking [1, 2, 3] revealed that black holes demonstrate thermodynamic properties, radiating thermal energy at a specific temperature—what we now call Hawking radiation. Nevertheless, this Hawking temperature remains extraordinarily low for astronomical black holes, creating significant obstacles for direct experimental observation. Additionally, [3] overlooked a crucial element—the backreaction, meaning the influence of quantum fluctuations on the background spacetime. Since the direct investigation of such systems is not possible, these constraints can rather be addressed and investigated in laboratory settings by developing quantum simulators for such phenomena using ultracold atomic gases [4, 5]. In this context, Bose-Einstein condensates (BECs) emerge as particularly valuable tools, offering precise experimental control alongside well-developed microscopic theoretical frameworks [6]. The quantum simulation of such spacetime using superfluids originated from Unruh’s proposal in 1981 [7], in which he showed that transonic fluid flows can mimic event horizons and suggested that black hole evaporation processes might be observable in such systems. Since then, BECs have emerged as a highly favorable platform, with experimental observations of analog Hawking radiation reported in [8, 9, 10, 11], reinforcing the feasibility of using quantum gases as simulators for such spacetimes.
While other platforms, such as superfluid helium, nonlinear optics, hydrodynamical analogs, and superconducting circuits [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], have been explored, BECs remain particularly compelling [25]. Numerous theoretical and experimental studies have since examined sonic black holes in BECs [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], as well as early universe analogs [49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Notably, ring-shaped (toroidal) condensates [59, 60, 61, 62, 27, 63] have gained attention for their intrinsic stability and periodic boundary conditions, offering a versatile setup for analog gravity simulations.
Alongside reproducing the kinematic aspects of curved spacetime, BEC-based quantum simulators go beyond and enable the controlled study of microscopic backreaction effects, otherwise inaccessible in real gravitational systems [64, 65]. Despite the intuitive assumption that Bogoliubov theory could easily handle backreaction in dilute BECs, this turns out to be nontrivial [65]. Though the broader research goal would be to investigate the correlation between Hawking temperature and backreaction [64, 66, 67, 65, 68, 69, 70] in such systems, via this work, however, we take a crucial fundamental step by examining the correlation between Hawking temperature $T_{H}$ and quantum depletion—the fraction of atoms that exit the condensate due to interactions and dynamic evolution.
Figure 1: Artist’s impression of ring BEC featuring acoustic black hole (left) and white hole (right) horizons. The state at $t=0$ corresponds to an initial configuration with no quantum depletion. As time progresses to $t=t_{m}$, quantum depletion begins to evolve, eventually reaching a maximal value at $t=t_{f}$, beyond which the background can no longer be considered a condensate, with the quantum component dominating over the coherent background that characterized the initial state.
In this letter, we study a one-dimensional ring-shaped Bose-Einstein condensate hosting an analog black hole (BH) and white hole (WH) horizon pair [71], as depicted in Fig. 1. This geometry provides a controlled and stable background for simulating Hawking-like processes, allowing us to dynamically evolve the system using the time-dependent Gross-Pitaevskii equation (GPE) and analyze quantum fluctuations via the Bogoliubov–de Gennes (BdG) formalism. The initial state is a fully condensed vacuum with no depletion, evolving in time as interactions and Hawking radiation-induced excitations generate depletion.
Our primary aim is to determine the threshold of $T_{H}$ beyond which the condensate ceases to be well-described by Bogoliubov theory—a limit corresponding to significant quantum depletion and the onset of strong backreaction. We find that quantum depletion increases monotonically with $T_{H}$, and that once a critical level is reached, the coherence of the condensate is lost, giving way to a regime dominated by quantum fluctuations.
Construction of black hole-white hole pair on ring. We consider a non-relativistic dilute bosonic gas, described by the field $\phi=\phi(t,x)$ in $1+1D$, which on the mean-field level is described by the GPE $(\hbar=1)$
$i\partial_{t}\phi=-\frac{1}{2m}\partial^{2}_{x}\phi+(V+g|\phi|^{2})\phi.$ (1)
with $m$ the boson mass, $g$ the coupling strength, and $V$ representing the external potential. We prepare the condensate to be in a ring-like configuration of total length $l$, satisfying periodic boundary conditions $\phi(t,-l/2)=\phi(t,l/2)$ and $\partial_{x}\phi(t,x)|_{x=-l/2}=\partial_{x}\phi(t,x)|_{x=l/2}$. We assume that for $t<0$, the condensate density, $\rho_{0}$ is constant, $g=0$, and then switch to a $g_{0}>0$. Accordingly, using (1),
$\frac{1}{g_{0}\rho_{0}}i\partial_{t}\phi=-\frac{\xi_{0}^{2}}{2}\partial_{x}^{2}\phi+\bigg{(}\frac{V}{g_{0}\rho_{0}}+\frac{g|\phi|^{2}}{g_{0}\rho_{0}}\bigg{)}\phi,$ (2)
where we defined $\xi_{0}=1/\sqrt{mg_{0}\rho_{0}}$. In what follows, we use as the unit of length $\xi_{0}$, thus time in units of $\xi_{0}^{2}$, $g$ is in units $g_{0}$ and $\phi$ in units of $\sqrt{\rho_{0}}$. Then we have
$i\partial_{t}\phi=-\frac{1}{2}\partial^{2}_{x}\phi+(V+g|\phi|^{2})\phi,$ (3)
[…]
$g(x)=1-\frac{e^{-x^{2}}}{2}.$ (4)
and is turned on at $t>0$, which completes our setup for constructing the black hole-white hole (BH-WH) pair in our ring condensate.
Horizons and background evolution. To isolate the effect of the horizons on quantum depletion, we explore our basic setup’s dynamical response in two cases: The BH-WH pair is present or it is not. Having fixed $V$ and $g(x)$, the local Mach number ${\mathfrak{M}}=v/c$, with $v$ and $c$ being flow velocity and sound speed, respectively, is
${\mathfrak{M}}=\frac{k_{n}}{\sqrt{g|\phi|^{2}}}$ (5)
at $t=0$, with ${\mathfrak{M}}=1$ representing the location of analogue event horizons. The presence or absence of the acoustic horizons can be controlled via the parameter $n$ in $k_{n}=\frac{2n\pi}{l}$, considering that the background density i.e. $|\phi|^{2}$ and $l$ remains fixed. The length of the ring $(l)$ is taken to be 20. Mach number and interaction $g(x)$ are shown in Figs.2. When we consider $n=3$, it establishes the BH-WH horizons at the red dot positions that represent the points on the ring where the Mach number reaches unity coming from below (BH) and above (WH). Horizons are absent when substituting $n=2$, cf. inset of Fig. 2.
Figure 2: mach number (orange) and g(x) (blue) vs x for n=3. The black line represents $y=1$ and is used for reference. The red dot represents the points on the x-axis where the Mach number crosses one, hence implying the presence of acoustic black hole and white hole horizons.
Figure 3: Background condensate for $n=2$ (blue) when there are no horizons, and for $n=3$ (Red) when horizons are present.
We now proceed to investigate the dynamics of the condensate for both scenarios, then adding quantum depletion dynamics later on. For $t>0$, we take the ansatz $\phi(t,x)=e^{-i\mu t+ik_{n}x}\varphi$, where $\varphi$ is a solution of
$i\partial_{t}\varphi=-\frac{1}{2}\partial^{2}_{x}\varphi-ik_{n}\partial_{x}\varphi+g|\phi|^{2}\phi,$ (6)
subject to the initial condition $\varphi(0,x)=1$.To study the evolution of background, we solve the equation 6 numerically when the BH-WH pair is present and when it is not. We show the results obtained for $|\phi|^{2}$ on solving (6), in Fig. 3 We conclude from Fig. 3 that $|\phi|^{2}$ displays a similar behavior for both cases. Here, we ran the simulation until $t=3.9$; our background is rendered stable in the time window $[0,3.9]$; $\varphi(0,x)=1$ was taken to be the initial solution.
Quantum depletion. Introducing a small fluctuating field $\chi$ on top of the background condensate $\phi$ as $\phi\rightarrow\phi+\chi$, using 1, we obtain the following Bogoliubov-de Gennes (BdG) equation for $\chi$
$i\partial_{t}\chi=-\frac{1}{2}\partial^{2}_{x}\chi+(V+2g|\phi|^{2})\chi+g\phi^{2}\chi^{*}.$ (7)
Here, we let $\chi=exp(-i\mu t+ik_{n}x)\psi$, such that
$i\partial_{t}\psi=-\frac{1}{2}\partial^{2}_{x}\psi-ik_{n}\partial_{x}\psi+2g|\varphi|^{2}\psi+g\varphi^{2}\psi^{*}.$ (8)
We define the density of the depleted cloud as $\rho_{\chi}=\langle\hat{\chi}^{\dagger}\hat{\chi}\rangle$. Using the Nambu spinor $\Psi$, defined as $\Psi=\begin{pmatrix}\psi\\ \psi^{*}\end{pmatrix}$, such that (7) can be written as
$i\partial_{t}\Psi=-\frac{1}{2}\partial^{2}_{x}\Psi-ik_{n}\sigma_{3}\partial_{x}\Psi+2g|\varphi|^{2}\Psi+g\sigma_{\varphi}\Psi^{*},$ (9)
where $\sigma_{\varphi}=\begin{pmatrix}0&\varphi^{2}\\ \varphi^{*2}&0\end{pmatrix}.$ For $t<0$, such a set of modes is $\Psi=\frac{1}{\sqrt{l}}e^{-i\omega_{j}t+ik_{j}x}\begin{pmatrix}1\\ 0\end{pmatrix}$ with $j$ any integer, and the frequencies $\omega_{j}=\frac{k_{j}^{2}}{2}+k_{j}k_{n}.$Here, $\Psi_{j}$ will evolve as the interactions are turned on ($t>0$). To calculate $\rho_{\chi}$, we proceed to solve Eq. 9 numerically with the initial condition $\Psi(0,x)=\frac{1}{\sqrt{l}}e^{ik_{j}x}\begin{pmatrix}1\\ 0\end{pmatrix}$ We then define the global quantum depletion $\mathfrak{D}$ as
[…]
and calculate it by evaluating $\rho_{\chi}$ via the solution of (9), and hence of (7). The results obtained are shown in Fig. 4. Again the red curve represents the evolution of depletion when a BH–WH pair is present in the condensate, while blue is without. This comparison clearly indicates that the presence of horizons affects depletion: We obtain a more depleted cloud when horizons are present. We now proceed explore how increasing the final Hawking temperature affects the quantum depletion.
Figure 4: Global quantum depletion $(\mathfrak{D})$ as a function of time with (red) and without (blue) horizons.
Figure 5: Evolution of the condensate density at different times for $T_{H}=0.027mc_{0}^{2}$.
Hawking temperature limit. The previous section demonstrated that the presence of the horizons impacts the depletion magnitude. We now relate the latter to the Hawking temperature. We also aim in this connection to determine the intimately related conditions under which Bogoliubov-de Gennes (BdG) theory remains valid. The analog Hawking temperature $(T_{H})$, which in a quasistationary limit is obtained in the horizonful configuration to which we quench [72, 73] $(k_{B}=1)$
$T_{H}=-\frac{\hbar}{2\pi}\frac{1}{2v_{h}}\left.\frac{d}{dx}(c^{2}-v^{2})\right|_{x=x_{H}}$ (11)
For our paramters $\rho_{0}=1$, $l=20$, we find $T_{H}\sim 0.027$ in units of $mc_{0}^{2}=1/\xi_{0}^{2}$. For $T_{H}\sim 0.027mc_{0}^{2}$, the quantum global depletion evolution is shown in Fig. 4. It reaches a maximum value of $0.10$ for this value of $T_{H}$. The background evolution until $t=10$ is shown in Fig. 5. This temperature is significantly lower than experimentally reported ones [8, 9, 10, 11]. To more closely approach these, we gradually increase $T_{H}$ via changing $\rho_{0}$ by up to a factor of ten, maintaining the acoustic horizon pair within the ring-condensate. The results is presented in Fig. 6, which shows the depletion as a function of both, $T_{H}$ and time after quenching to the BH-WH pair, $t$. As $T_{H}$ increases, the depletion also increases—even when the condensate run time is held constant across different temperatures—indicating a clear dependence of depletion on the Hawking temperature. We increase $T_{H}$ up to $0.1$ in our unit $mc_{0}^{2}$, at which point the depletion is observed to be approximately $25\%$, cf. Fig. 6. At this level of depletion, around $75\%$ of the particles still occupy the condensate mode. However, the depletion is no longer negligible, and beyond-mean-field effects begin to play a significant role. Any further increase in $T_{H}$ leads to an increasingly rapid growth of depletion, and ultimately a macroscopic occupation of the condensate mode no longer exists. Accordingly, the Bogoliubov-de Gennes (BdG) equation then is rendered invalid, and a many-body treatment taking into account strong correlations is then required. The accelerated depletion on further increase of $T_{H}$ not only pushes the system closer to the breakdown of condensation but also significantly enhances backreaction effects, and quantum excitations begin to influence the background dynamics non-negligibly. The magnitude of fluctuations becomes comparable to that of the condensate itself, necessitating the fully quantum many-body or stochastic field-theoretic treatment. However, a notable advantage of observing depletion at higher values of $T_{H}$—still remaining below the upper threshold of this study i.e. $T_{H}\sim 0.1mc_{0}^{2}$, is that experimental detection of quantum depletion becomes more feasible. The depletion increases rapidly, making it easier to distinguish it from background noise or from finite-temperature effects within the experimental runtime. We also explored an alternative method to increase $T_{H}$ by spatial variation of the interaction strength $g(x)$ in (4). This has led, however, to the horizons moving too close to one another, thus creating a “quasi-microscopic” BH-WH pair and making it difficult to resolve Hawking-like effects or local features of the quantum depletion. Modulating the initial condensate density proved to be a more effective and controllable strategy for increasing $T_{H}$ while preserving the global structure around the ring, allowing for a clearer observation of the depletion dynamics with time.
Figure 6: Time evolution of global quantum depletion $(\mathfrak{D})$ for various values of $T_{H}$. The color bar on the right represents the corresponding depletion scale.
Conclusion and outlook. In this work, we investigated the correlation between the Hawking temperature for a ring-shaped Bose-Einstein condensate (BEC) hosting an analog black hole–white hole (BH-WH) pair, and quantum depletion as the system evolves, following the Bogoliubov approximation. By comparing the depletion dynamics in the presence and absence of horizons, we demonstrated that the BH–WH configuration leads to a relatively faster growth in depletion, indicating a clear interplay between horizon-induced phonon production and condensate coherence loss. The primary goal of the present study was to identify regimes of large Hawking temperature $T_{H}$ while still being able to maintain the validity of the Bogoliubov approximation of excitations propagating above the condensate. Using density as a tunable parameter, we explored a range of $T_{H}$ values and determined that a reliable analysis of global quantum depletion is possible up to $T_{H}\sim 0.1mc_{0}^{2}$. At this point, the depletion reaches around $25\%$, implying that the system still retains a condensate fraction of roughly $75\%$. Beyond this point, quantum fluctuations become substantial and beyond-mean-field effects begin to emerge. We found that further increase in $T_{H}$ leads to accelerated depletion, rapid evolution of the background, and onset of strong backreaction effects, which challenge the validity of the BdG framework. Once depletion becomes comparable to the condensate density, the system approaches the breakdown of condensation, and the Gross-Pitaevskii plus Bogoliubov description becomes insufficient. Therefore, the regime beyond $T_{H}\sim 0.1mc_{0}^{2}$ requires a fully many-body treatment or inclusion of stochastic quantum corrections.
In terms of methods for controlling $T_{H}$, modulating the interaction strength $g(x)$ was also thought of but, we found that it causes undesirable shifts in the horizon positions—compressing the BH–WH region and limiting observational resolution—while the density variation provided a more controlled means of increasing $T_{H}$ while preserving global structure, making it better suited for experimental and theoretical investigations of quantum depletion dynamics.
Finally, our results highlight a key trade-off: Although higher $T_{H}$ values allow for faster observation of depletion effects—potentially within accessible experimental timescales—they also bring the system closer to regimes where strong backreaction and breakdown of the condensate set in. This work takes a fundamental step towards setting up a tabletop experiment to investigate the missing factor in [2, 3], i.e., otherwise inaccessible in direct observation. Future studies will address going beyond the Bogoliubov approximation and the impact of strong quantum backreaction on analogue Hawking radiation. Additionally, studying the microscopic structure and evolution of depletion near the BH–WH pair may offer deeper insights into horizon-induced quantum effects.
Acknowledgment. AR gratefully acknowledges Prof. Uwe R. Fischer for conceiving the original idea for this project and for his continuous guidance and insightful contributions throughout all stages of this work, including detailed feedback on the manuscript. Special thanks are extended to Dr. Caio C. Holanda Ribeiro for his input on the development of the initial setup and for many valuable discussions that significantly shaped the progress of the project. This work was supported by the National Research Foundation of Korea under Grant No. 2020R1A2C2008103.